# Sifan Wang Graduate Group in Applied Mathematics and Computational Science

### 01 Feb 2024 v1

@@ 1 Introduction
Machine learning (ML) is making continual impact to the fields of science and engineering, providing advanced tools for analyzing complex data, uncovering nonlinear relationships and developing predictive models. Notable examples include AlphaFold [1] in protein structure prediction, Deep Potentials [2] for large-scale molecular dynamics, and GraphCast [3] for medium-range weather forecasting. The integration of physical laws and constraints within machine learning models has given rise to physics-informed machine learning (PIML). This emerging field opens new frontiers for traditional scientific research, and contributes to address persistent challenges in machine learning, such as robustness, interpretability, and generalization.
@@
The fundamental question that PIML aims to address is how to incorporate physical prior knowledge into machine learning models. This can be achieved by modifying key components of the machine learning pipeline, which includes data processing, model architecture, loss functions, optimization algorithms, and fine-tuning and inference. For example, neural operators [4, 5] are designed to encode physics from training data that reflects underlying physical laws. Another powerful approach is to tweak network architectures to strictly enforce physical constraints, thereby enhancing the generalizability of the model. One method involves embedding general physical principles such as symmetries, invariances, and equivariances into ML models. As highlighted by Cohen et al. [6] and Maron et al. [7, 8], this strategy leads to simpler models that require less data while achieving higher prediction accuracy. Alternatively, embedding specific physical laws can be equally effective. This is demonstrated by Mohan et al.[9], who ensured continuity and mass conservation in neural networks for coarse-graining three-dimensional turbulence by integrating the curl operator. Similarly, Meng et al.[10] illustrate that certain neural network architectures are inherently aligned with the physics of specific Hamilton-Jacobi (HJ) partial differential equations (PDEs).
@@
One of the most popular and flexible methods for embedding physical principles in machine learning is through the formulation of tailored loss functions. These loss functions serve as soft constraints that bias ML models towards respecting the underlying physics during training, giving rise to the emergence of physics-informed neural networks (PINNs). Thanks to their flexibility and ease of implementation, PINNs have been extensively used to solve forward and inverse problems involving PDEs by seamlessly integrate noisy experimental data and physical laws into the learning process. In recent years, PINNs have yielded a series of promising results across various domains in computational science, including applications in fluid mechanics [11, 12, 13], bio-engineering [14, 15], and material science [16, 17, 18]. Furthermore, PINNs have been effectively applied in molecular dynamics [19], electromagnetics [20, 21], geosciences [22, 23], and in designing thermal systems [24, 25].
@@
Despite significant progress, most existing work on PINNs tends to employ small, and shallow network architectures, leaving the vast potential of deep networks largely untapped. To bridge this gap, we put forth a novel class of architectures coined Physics-Informed residual adaptive networks (PirateNets). Our main contributions are summarized as follows:
@@
  • We empirically and theoretically reveal that conventional initialization schemes (e.g., Glorot or He initialization) result in problematic initialization of MLP derivatives and thus worse trainablity.
@@
  • We conduct comprehensive numerical experiments demonstrating that PirateNets achieve consistent improvements in accuracy, robustness and stability across various benchmarks.
The remaining of this paper is organized as follows. Section 2 offers a concise overview of the original formulation of PINNs, as introduced by Raissi et al. [54]. In Section 3, we start with an example that illustrates the growing instability in training PINNs with deep neural networks. We then put forth a hypothesis: if a PINN model is trained sufficiently, then both its predictions and their derivatives should closely approximate the solution of the underlying PDE and its derivatives. Further, we prove that this hypothesis holds for linear elliptic and parabolic PDEs. Based on this assumption, in Section 3, we investigate the capabilities of MLP derivatives, uncovering that deep MLP derivatives often suffer from ill-posed initializations, which contribute to an unstable training process. To address this issue, we propose PirateNets in Section 4. In Section 5, we perform comprehensive numerical experiments to validate the proposed architecture. Finally, Section 6 concludes the paper, summarizing our key findings and discussing potential future research directions.
@@ 2 Physics-informed neural network
We begin with a brief overview of physics-informed neural networks (PINNs) following the original formulation of Raissi et al. [54]. Generally, we consider solving a PDE of the form
@@
where $\mathcal{N}[\cdot]$ is a linear or nonlinear differential operator and $\mathbf{u}(t,\mathbf{x})$ describes the unknown solution defined in a spatio-temporal domain with $t\in[0,T]$ and $\mathbf{x}\in\Omega\subset\mathbb{R}^{d}$. The initial and boundary conditions are given by
$\mathbf{u}(0,\mathbf{x})=\mathbf{g}(\mathbf{x}),\ \ \mathbf{x}\in\Omega,$ (2.2)
$\mathcal{B}[\mathbf{u}]=0,\ \ t\in[0,T],\ \mathbf{x}\in\partial\Omega.$ (2.3)
Here, $\mathcal{B}[\cdot]$ denotes a general boundary operator, which can represent various boundary conditions, including Dirichlet, Neumann, Robin, periodic conditions, etc..
We proceed by approximating the unknown solution $\mathbf{u}(t,\mathbf{x})$ by a deep neural network $\mathbf{u}_{\mathbf{\theta}}(t,\mathbf{x})$, where $\mathbf{\theta}$ denotes all trainable parameters of the network (e.g., weights and biases). This allows us to define the PDE residuals as
$\mathcal{R}[u_{\mathbf{\theta}}](t,\mathbf{x})=\frac{\partial\mathbf{u}_{\mathbf{\theta}}}{\partial t}(t,\mathbf{x})+\mathcal{D}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x}).$ (2.4)
Note that $\mathbf{u}_{\mathbf{\theta}}$ provides a continuous representation that can be queried for any $(t,\mathbf{x})$. As a result, all required gradients with respect to input variables or network parameters $\mathbf{\theta}$ can be computed via automatic differentiation [55]. Then, we can train a physics-informed model by minimizing the following composite loss
$\mathcal{L}(\mathbf{\theta})=\mathcal{L}_{ic}(\mathbf{\theta})+\mathcal{L}_{bc}(\mathbf{\theta})+\mathcal{L}_{r}(\mathbf{\theta}),$ (2.5)
where
$\mathcal{L}_{ic}(\mathbf{\theta})=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left|\mathbf{u}_{\mathbf{\theta}}(0,\mathbf{x}_{ic}^{i})-\mathbf{g}(\mathbf{x}_{ic}^{i})\right|^{2},$ (2.6)
$\mathcal{L}_{bc}(\mathbf{\theta})=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left|\mathcal{B}[\mathbf{u}_{\mathbf{\theta}}](t_{bc}^{i},\mathbf{x}_{bc}^{i})\right|^{2},$ (2.7)
$\mathcal{L}_{r}(\mathbf{\theta})=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left|\mathcal{R}[u_{\mathbf{\theta}}](t^{i}_{r},\mathbf{x}^{i}_{r})\right|^{2}.$ (2.8)
These loss functions aim to enforce the network to fit the initial, boundary conditions and to satisfy the PDEs, respectively. Besides, $\{\mathbf{x}_{ic}^{i}\}_{i=1}^{N_{ic}}$, $\{t_{bc}^{i},\mathbf{x}_{bc}^{i}\}_{i=1}^{N_{bc}}$ and $\{t_{r}^{i},\mathbf{x}_{r}^{i}\}_{i=1}^{N_{r}}$ can be the vertices of a fixed mesh or points that are randomly sampled at each iteration of a gradient descent algorithm.
@@ 3 Initialization pathologies in PINNs
Training PINNs is notably more difficult than supervised training of conventional neural networks. As demonstrated by Wang et al., PINNs face several training pathologies including spectral bias [26, 27], unbalanced losses [28], and causality violation [30]. Efforts to mitigate these issues include the use of Fourier feature networks [56], self-adaptive weighting schemes [28, 29] and the causal and curriculum training algorithms [30, 49, 50]. However, despite these advances, the practical deployment of PINNs has been largely limited to employing small and shallow networks, typically with 5 layers or less. This is in sharp contrast to the exceptional capabilities showcased by deeper networks in the broader field of deep learning [57, 58, 59].
In the following, we will both empirically and theoretically demonstrate that the trainability of PINNs degrades as the network depth increases, particularly when employing multi-layer perceptrons (MLPs) as the backbone architecture. To this end, we consider the Allen-Cahn equation, a challenging benchmark for conventional PINN models that has been extensively studied in recent literature [49, 30, 60, 34, 61]. The equation is defined as follows,
$u_{t}-0.0001u_{xx}+5u^{3}-5u=0,\quad t\in[0,1],x\in[-1,1],$ (3.1)
$u(0,x)=x^{2}\cos(\pi x),$ (3.2)
$u(t,-1)=u(t,1),$ (3.3)
$u_{x}(t,-1)=u_{x}(t,1).$ (3.4)
@@
Figure 1: Allen-Cahn equation: Relative $L^{2}$ error of training PINNs using MLP, ResNet and PirateNet backbones of varying depth, averaged over 5 random seeds for each architecture.
Figure 1 shows the resulting relative $L^{2}$ errors when training PINNs using MLPs of varying depth, averaged over 5 different random seeds. It can be observed that the prediction error becomes larger as the network depth increases. Eventually, this error almost reaches $100$% for PINN models with 18 hidden layers, indicating a complete failure of the model to yield a reasonable solution.
One may argue that the challenges in training deep neural networks can be effectively resolved by incorporating skip connections [64]. To test this hypothesis, our study compares the performance of MLP networks with residual connections (ResNets) under identical hyper-parameter settings. As depicted in the same figure, the overall trend remains consistent: the error proportionally increases with network depth, albeit at a slightly slower rate. This implies that, contrary to expectations, the conventional skip connection fails to address the fundamental issues in training PINNs using deeper networks. These findings highlight a critical limitation in PINN performance when the network depth exceeds a certain threshold. To address this challenge, To address this challenge, we propose a novel network architecture in Section 4, referred to as Physics-Informed Residual AdapTivE Networks (PirateNets), designed for efficient and stable training of deep PINN models. As can be seen in Figure 1, PirateNets consistently reduces the predictive error as the network depth increases.
To gain deeper insights into the reasons behind the poor performance of MLPs and ResNets, we focus on examining the initialization process of PINNs. Because of the variety and nonlinearity in PDE expressions, a direct analysis of the initialization of the PDE residual network (denoted by $\mathcal{R}[\mathbf{u}_{\mathbf{\theta}}]$) presents significant challenges. To simplify this complexity, let us revisit the role of the PDE residual loss 2.8, which aims to enforce the network to satisfy the underlying PDEs. From this perspective, it is reasonable to expect that, if the training loss is small, then the network’s derivatives should closely approximate those of the actual solution. More formally, we put forth the following assumption:
@@ 4 Physics-informed residual adaptive networks (PirateNets)
In this section, we introduce Physics-Informed Residual AdapTivE Networks (PirateNets), a novel architecture designed to address the issue demonstrated in the previous sections. Figure 3 illustrates the forward pass of a PirateNet. First, the input coordinates $\mathbf{x}$ are transformed into a high dimensional feature space using an embedding function $\Phi(\mathbf{x})$. In this work we employ random Fourier features [56] as the the embedding. This method has been extensively validated for reducing spectral bias in the training of PINNs, thus enabling more efficient approximation of high-frequency solutions [27]. The embedding $\Phi(\mathbf{x})$ is defined as:
$\Phi(\mathbf{x})=\begin{bmatrix}\cos(\mathbf{Bx})\\ \sin(\mathbf{Bx})\end{bmatrix},$ (4.3)
where each entry in $\mathbf{B}\in\mathbb{R}^{m\times d}$ is sampled from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ and $\sigma>0$ is a user-specified hyper-parameter.
The embedded coordinates $\Phi(\mathbf{x})$ are then mapped into two dense layers
$\mathbf{U}$ $=\sigma(\mathbf{W}_{1}\Phi(\mathbf{x})+\mathbf{b}_{1}),$ (4.4)
$\mathbf{V}$ $=\sigma(\mathbf{W}_{2}\Phi(\mathbf{x})+\mathbf{b}_{2}),$ (4.5)
where $\sigma$ denotes a point-wise activation function. These two encoding maps act as gates in each residual block of the architecture. This choice is motivated by [28] and has been widely used to enhance the trainability of MLPs and improve the convergence of PINNs [62, 61].
Let $\mathbf{x}^{(l)}$ denote the input of the $l$-th block, respectively. The forward pass within each PirateNet block is defined as
$\mathbf{f}^{(l)}$ $=\sigma(\mathbf{W}^{(l)}_{1}\mathbf{x}^{(l)}+\mathbf{b}^{(l)}_{1}),$ (4.6)
@@
$\mathbf{g}^{(l)}$ $=\sigma(\mathbf{W}^{(l)}_{2}\mathbf{z}_{1}^{(l)}+\mathbf{b}^{(l)}_{2}),$ (4.8)
@@
$\mathbf{h}^{(l)}$ $=\sigma(\mathbf{W}^{(l)}_{3}\mathbf{z}_{2}^{(l)}+\mathbf{b}^{(l)}_{3}),$ (4.10)
$\mathbf{x}^{(l+1)}$ $=\alpha^{(l)}\cdot\mathbf{h}^{(l)}+(1-\alpha^{(l)})\cdot\mathbf{x}^{(l)},$ (4.11)
where $\odot$ denotes a point-wise multiplication and $\alpha^{(l)}\in\mathbb{R}$ is a trainable parameter. All the weights are initialized by the Glorot scheme [67] and biases are initialized to zero.
@@
Notice that each residual block consists of three dense layers with two gating operations, followed by an adaptive residual connection across the stacked layers. It is worth noting that a PirateNet of $L$ residual blocks will have depth of $3L$ and its total number of trainable parameters is comparable to that of a MLP of the same depth.
A key aspect of PirateNets is the trainable parameters $\alpha^{(l)}$ in the skip connections, which determine the nonlinearity of the $l$-th block. Specifically, when $\alpha^{(l)}=0$, $\mathbf{x}^{(l+1)}=\mathbf{x}^{(l)}$ is an identity map. Conversely, when $\alpha^{(l)}=1$, the mapping becomes fully nonlinear without any shortcuts. Throughout all experiments in this work, we initialize $\alpha^{(l)}$ to zero for all blocks, leading to the final output of a PirateNet $\mathbf{u}_{\mathbf{\theta}}$ being a linear combination of the first layer embeddings at initialization, i.e.,
@@
By doing so, we circumvent the initialization pathologies of deep networks identified in Section 3. PirateNets are designed to learn the requisite nonlinearity, $\alpha$, based on the PDE system during the training process. As a result, both their trainability and expressivity are restored, with the models’ forward pass becoming more nonlinear and deeper. In fact, the PDE solution can be simply approximated by a small shallow network or even a linear combination of a basis, similar to spectral and finite element methods. The rationale behind employing deep neural networks lies in leveraging additional nonlinearities to minimize the PDE residuals, thereby enabling the network’s derivatives to learn the solution and its derivatives accordingly.
From Equation 4.13, another key observation is that PirateNets can be viewed as a linear combination of basis functions at initialization. This not only allows one to control the inductive bias of the network by appropriate choice of basis, but it also enables the integration of various types of existing data into the network’s initialization phase. Specifically, suppose that there is a set of solution measurements, denoted as $\mathbf{Y}=\{y_{i}\}_{i=1}^{n}$. These can be readily used to appropriately initialize the last linear layer of the model by solving a linear least square problem,
@@
Figure 3: Physics-informed residual adaptive networks (PirateNets): In our model, input coordinates are first projected into high-dimensional feature spaces using random Fourier features, then followed by passing through $N$ adaptive residual blocks. Each block consists of three dense layers, augmented with two gating operations that incorporate shallow latent features. The key design of the architecture is the adaptive skip connection with a trainable parameter $\alpha$ initialized at 0. This ensures an identity mapping for each block, thereby initializing the model as a linear combination of the coordinate embeddings. More importantly, this approach circumvents the issue of pathological initialization in deep PDE residual networks. For the final layer, we propose a physics-informed initialization based on solving a least squares problem to fit the available data, while all other weights are initialized following the Glorot scheme, and biases are set to zero. As training progresses, the model’s depth gradually increases as the nonlinearities become activated, enabling the model to progressively recover its approximation capacity.

### 05 Feb 2024 v2

@@ 1 Introduction
Machine learning (ML) is making a continual impact on the fields of science and engineering, providing advanced tools for analyzing complex data, uncovering nonlinear relationships, and developing predictive models. Notable examples include AlphaFold [1] in protein structure prediction, Deep Potentials [2] for large-scale molecular dynamics, and GraphCast [3] for medium-range weather forecasting. The integration of physical laws and constraints within machine learning models has given rise to physics-informed machine learning (PIML). This emerging field opens new frontiers for traditional scientific research and contributes to addressing persistent challenges in machine learning, such as robustness, interpretability, and generalization.
@@
The fundamental question that PIML aims to address is how to incorporate physical prior knowledge into machine learning models. This can be achieved by modifying key components of the machine learning pipeline, which includes data processing, model architecture, loss functions, optimization algorithms, and fine-tuning and inference. For example, neural operators [4, 5] are designed to encode physics from training data that reflects underlying physical laws. Another powerful approach is to tweak network architectures to strictly enforce physical constraints, thereby enhancing the generalizability of the model. One method involves embedding general physical principles such as symmetries, invariances and equivariances into ML models. As highlighted by Cohen et al. [6] and Maron et al. [7, 8], this strategy leads to simpler models that require less data while achieving higher prediction accuracy. Alternatively, embedding specific physical laws can be equally effective. This is demonstrated by Mohan et al.[9], who ensured continuity and mass conservation in neural networks for coarse-graining three-dimensional turbulence by integrating the curl operator. Similarly, Meng et al.[10] illustrates that certain neural network architectures are inherently aligned with the physics of specific Hamilton-Jacobi (HJ) partial differential equations (PDEs).
@@
One of the most popular and flexible methods for embedding physical principles in machine learning is through the formulation of tailored loss functions. These loss functions serve as soft constraints that bias ML models towards respecting the underlying physics during training, giving rise to the emergence of physics-informed neural networks (PINNs). Thanks to their flexibility and ease of implementation, PINNs have been extensively used to solve forward and inverse problems involving PDEs by seamlessly integrating noisy experimental data and physical laws into the learning process. In recent years, PINNs have yielded a series of promising results across various domains in computational science, including applications in fluid mechanics [11, 12, 13], bio-engineering [14, 15], and material science [16, 17, 18]. Furthermore, PINNs have been effectively applied in molecular dynamics [19], electromagnetics [20, 21], geosciences [22, 23], and in designing thermal systems [24, 25].
@@
Despite significant progress in recent years, most existing works on PINNs tend to employ small, and shallow network architectures, leaving the vast potential of deep networks largely untapped. To bridge this gap, we put forth a novel class of architectures coined Physics-Informed residual adaptive networks (PirateNets). Our main contributions are summarized as follows:
@@
  • We empirically and theoretically reveal that conventional initialization schemes (e.g., Glorot or He initialization) result in problematic initialization of MLP derivatives and thus worse trainability.
@@
  • We conduct comprehensive numerical experiments demonstrating that PirateNets achieve consistent improvements in accuracy, robustness, and scalability across various benchmarks.
The rest of this paper is organized as follows. We first give a concise overview of physics-informed neural networks in Section 2, following the original formulation of Raissi et al. [54]. In Section 3, we take the Allen-Cahn equation as a motivating example to illustrate the growing instability in training PINNs with deep neural networks and then propose a potential rationale (mechanism): if a PINN model is trained sufficiently, then both its predictions and derivatives should well approximate the solution of the underlying PDE and its derivatives, respectively. We theoretically justify this claim for linear elliptic and parabolic PDEs and, by investigating the capabilities of MLP derivatives, find that deep MLP derivatives often suffer from ill-posed initializations, which could result in an unstable training process. To address the aforementioned issues, in Section 4 we propose a novel framework of PirateNets. In Section 5, comprehensive numerical experiments are performed to validate the proposed architecture. In Section 6, we conclude this work with a discussion of potential future research directions.
@@ 2 Physics-informed neural network
In this section, we will briefly overview the standard formulation of PINNs, following the one in [54]. Without loss of generality, we consider an abstract PDE of the form:
@@
defined on a spatial-temporal domain $[0,T]\times\Omega\subset\mathbb{R}^{1+d}$, where $\Omega$ is a bounded domain in $\mathbb{R}^{d}$ with regular enough boundary $\partial\Omega$, $\mathcal{D}[\cdot]$ is a linear or nonlinear differential operator, and $\mathbf{u}(t,\mathbf{x})$ denotes a unknown solution. The general initial and boundary conditions can be then formulated as:
$\mathbf{u}(0,\mathbf{x})=\mathbf{g}(\mathbf{x}),\quad\mathbf{x}\in\Omega\,,$ (2.2)
$\mathcal{B}[\mathbf{u}]=0,\quad t\in[0,T],\ \mathbf{x}\in\partial\Omega\,,$ (2.3)
with $\mathbf{g}(\mathbf{x})$ being a given function with certain regularity. Here, we denote by $\mathcal{B}[\cdot]$ an abstract boundary operator, and (2.3) can represent various boundary conditions such as Dirichlet, Neumann, Robin, and periodic conditions.
We aim at approximating the unknown solution $\mathbf{u}(t,\mathbf{x})$ by a deep neural network $\mathbf{u}_{\mathbf{\theta}}(t,\mathbf{x})$, where $\mathbf{\theta}$ denotes the set of all trainable parameters of the network (e.g., weights and biases). If a smooth activation function is employed, $\mathbf{u}_{\mathbf{\theta}}$ provides a smooth representation that can be queried for any $(t,\mathbf{x})$. Moreover, all required gradients with respect to input variables or network parameters $\mathbf{\theta}$ can be computed via automatic differentiation [55]. This allows us to define the interior PDE residuals as
$\mathcal{R}_{{\rm int}}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x})=\frac{\partial\mathbf{u}_{\mathbf{\theta}}}{\partial t}(t,\mathbf{x})+\mathcal{D}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x})\,,\quad(t,\mathbf{x})\in[0,T]\times\Omega\,,$ (2.4)
and spatial and temporal boundary residuals, respectively, by
$\mathcal{R}_{{\rm bc}}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x})=\mathcal{B}[\mathbf{u}_{\theta}](t,\mathbf{x})\,,\quad(t,\mathbf{x})\in[0,T]\times\partial\Omega\,,$
and
$\mathcal{R}_{{\rm ic}}[\mathbf{u}_{\mathbf{\theta}}](\mathbf{x})=\mathbf{u}_{\theta}(0,\mathbf{x})-\mathbf{g}(\mathbf{x})\,,\quad\mathbf{x}\in\Omega\,.$
Then, we train a physics-informed model by minimizing the following composite loss:
$\mathcal{L}(\mathbf{\theta}):=\underbrace{\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left|\mathcal{R}_{{\rm ic}}[\mathbf{u}_{\mathbf{\theta}}](\mathbf{x}_{ic}^{i})\right|^{2}}_{\mathcal{L}_{ic}(\mathbf{\theta})}+\underbrace{\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left|\mathcal{R}_{{\rm bc}}[\mathbf{u}_{\mathbf{\theta}}](t_{bc}^{i},\mathbf{x}_{bc}^{i})\right|^{2}}_{\mathcal{L}_{bc}(\mathbf{\theta})}+\underbrace{\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left|\mathcal{R}_{\rm int}[\mathbf{u}_{\mathbf{\theta}}](t^{i}_{r},\mathbf{x}^{i}_{r})\right|^{2}}_{\mathcal{L}_{r}(\mathbf{\theta})}\,,$ (2.5)
which aims to enforce the neural network function $\mathbf{u}_{\theta}$ to satisfy the PDEs (2.1) with initial and spatial boundary conditions (2.2)–(2.3). Besides, $\{\mathbf{x}_{ic}^{i}\}_{i=1}^{N_{ic}}$, $\{t_{bc}^{i},\mathbf{x}_{bc}^{i}\}_{i=1}^{N_{bc}}$ and $\{t_{r}^{i},\mathbf{x}_{r}^{i}\}_{i=1}^{N_{r}}$ can be the vertices of a fixed mesh or points that are randomly sampled at each iteration of a gradient descent algorithm.
@@ 3 Initialization pathologies in PINNs
It has been known for some time that training PINNs is notably more difficult than the supervised training of conventional neural networks. As demonstrated by Wang et al., PINNs face several training pathologies including spectral bias [26, 27], unbalanced losses [28], and causality violation [30]. Efforts to mitigate these issues include the use of Fourier feature networks [56], self-adaptive weighting schemes [28, 29] and the causal and curriculum training algorithms [30, 49, 50]. However, despite these advances, the practical deployment of PINNs has been largely limited to employing small and shallow networks, typically with 5 layers or less. This is in sharp contrast to the exceptional capabilities showcased by deeper networks in the broader field of deep learning [57, 58, 59].
In the following, we will both empirically and theoretically demonstrate that the trainability of PINNs degrades as the network depth increases, particularly when employing multi-layer perceptrons (MLPs) as the backbone architecture. To this end, let us first focus on the Allen-Cahn equation, a challenging benchmark for conventional PINN models that has been extensively studied in recent literature [49, 30, 60, 34, 61]. For simplicity, we consider the one-dimensional case with a periodic boundary condition:
$u_{t}-0.0001u_{xx}+5u^{3}-5u=0\,,\quad t\in[0,1]\,,\ x\in[-1,1]\,,$
$u(0,x)=x^{2}\cos(\pi x)\,,$
$u(t,-1)=u(t,1)\,,\quad u_{x}(t,-1)=u_{x}(t,1)\,.$
@@
Figure 1: Allen-Cahn equation: Relative $L^{2}$ error of training PINNs using MLP, ResNet, and PirateNet backbones of varying depths, averaged over 5 random seeds for each architecture.
Figure 1 shows the resulting relative $L^{2}$ errors when training PINNs with MLPs of varying depths. The result is averaged over 5 different random seeds. It can be observed that the prediction error becomes larger as the network depth increases. Eventually, this error almost reaches $100$% for PINN models with 18 hidden layers, indicating a complete failure of the model to yield a reasonable solution.
One may argue that the challenges in training deep neural networks can be effectively resolved by incorporating skip connections [64]. To test this hypothesis, our study compares the performance of MLP networks with residual connections (ResNets) under identical hyper-parameter settings. As depicted in the same figure, the overall trend remains consistent: the error proportionally increases with network depth, albeit at a slightly slower rate. This implies that, contrary to expectations, the conventional skip connection fails to address the fundamental issues in training PINNs using deeper networks. These findings highlight a critical limitation in PINN performance when the network depth exceeds a certain threshold. To address this challenge, we propose a novel network architecture in Section 4, referred to as Physics-Informed Residual AdapTivE Networks (PirateNets), designed for efficient and stable training of deep PINN models. As can be seen in Figure 1, PirateNets consistently reduces the predictive error as the network depth increases.
To gain deeper insights into the reasons behind the poor performance of MLPs and ResNets, we next examine the initialization process of PINNs. Because of the variety and nonlinearity in PDE expressions, a direct analysis of the initialization of the PDE residual network (denoted by $\mathcal{R}[\mathbf{u}_{\mathbf{\theta}}]$) presents significant challenges. To simplify this complexity, let us revisit the role of the PDE residual loss in (2.5), which aims to enforce the network to satisfy the underlying PDEs. From this perspective, it is reasonable to expect that, if the training loss is small, then the network’s derivatives should closely approximate those of the actual solution. More formally, we put forth the following assumption:
@@ 4 Physics-informed residual adaptive networks (PirateNets)
In this section, we introduce Physics-Informed Residual AdapTivE Networks (PirateNets), a novel architecture designed to address the initialization issue elaborated in Section 3. Figure 3 below illustrates the main modules of the forward pass of a PirateNet. In detail, the input coordinates $\mathbf{x}$ are first mapped into a high dimensional feature space by an embedding function $\Phi(\mathbf{x})$. In this work, we employ random Fourier features [56]:
$\Phi(\mathbf{x})=\begin{bmatrix}\cos(\mathbf{Bx})\\ \sin(\mathbf{Bx})\end{bmatrix},$
where each entry in $\mathbf{B}\in\mathbb{R}^{m\times d}$ is i.i.d. sampled from a Gaussian $\mathcal{N}(0,s^{2})$ with the standard deviation $s>0$ being a user-specified hyper-parameter. Such an embedding has been extensively validated for reducing spectral bias in the training of PINNs, thus enabling a more efficient approximation of high-frequency solutions [27].
Then, the embedded coordinates $\Phi(\mathbf{x})$ are sent into two dense layers:
$\mathbf{U}$ $=\sigma(\mathbf{W}_{1}\Phi(\mathbf{x})+\mathbf{b}_{1})\,,\quad\mathbf{V}=\sigma(\mathbf{W}_{2}\Phi(\mathbf{x})+\mathbf{b}_{2})\,,$
where $\sigma$ denotes a point-wise activation function. These two encoding maps act as gates in each residual block of the architecture. This step is motivated by [28] and has been widely used to enhance the trainability of MLPs and improve the convergence of PINNs [62, 61].
Let $\mathbf{x}^{(l)}$ denote the input of the $l$-th block for $1\leq l\leq L$. The forward pass in each PirateNet block is defined by the following iterations:
$\mathbf{f}^{(l)}$ $=\sigma\big{(}\mathbf{W}^{(l)}_{1}\mathbf{x}^{(l)}+\mathbf{b}^{(l)}_{1}\big{)}\,,$ (4.1)
@@
$\mathbf{g}^{(l)}$ $=\sigma\big{(}\mathbf{W}^{(l)}_{2}\mathbf{z}_{1}^{(l)}+\mathbf{b}^{(l)}_{2}\big{)}\,,$ (4.3)
@@
$\mathbf{h}^{(l)}$ $=\sigma\big{(}\mathbf{W}^{(l)}_{3}\mathbf{z}_{2}^{(l)}+\mathbf{b}^{(l)}_{3}\big{)}\,,$ (4.5)
$\mathbf{x}^{(l+1)}$ $=\alpha^{(l)}\mathbf{h}^{(l)}+(1-\alpha^{(l)})\mathbf{x}^{(l)}\,,$ (4.6)
where $\odot$ denotes a point-wise multiplication and $\alpha^{(l)}\in\mathbb{R}$ is a trainable parameter. All the weights are initialized by the Glorot scheme [67], while biases are initialized to zero.
@@
It is worth noting from (4.1)–(4.6) that each residual block consists of three dense layers with two gating operations, followed by an adaptive residual connection across the stacked layers. Thus, a PirateNet of $L$ residual blocks will have a depth of $3L$ and its total number of trainable parameters is comparable to that of an MLP of the same depth.
A key aspect of PirateNets is the trainable parameters $\alpha^{(l)}$ in the skip connections, which determine the nonlinearity of the $l$-th block. Specifically, when $\alpha^{(l)}=0$, $\mathbf{x}^{(l+1)}=\mathbf{x}^{(l)}$ by (4.6), meaning that in this case, the $l$-th block is an identity map. In contrast, when $\alpha^{(l)}=1$, the mapping becomes fully nonlinear without any shortcuts. Throughout all experiments in this work, we initialize $\alpha^{(l)}$ to zero for all blocks, leading to the final output of a PirateNet $\mathbf{u}_{\mathbf{\theta}}$ being a linear combination of the first layer embeddings at initialization, i.e.,
@@
By doing so, we circumvent the initialization pathologies of deep networks discussed in Section 3. PirateNets are designed to learn the requisite nonlinearity encoded in $\alpha$, based on the PDE system, during the training process. As a result, both their trainability and expressivity are restored, with the models’ forward pass becoming more nonlinear and deeper. In fact, the PDE solution can be simply approximated by a small shallow network or even a linear combination of some basis, similar to spectral and finite element methods. The rationale behind employing deep neural networks lies in leveraging additional nonlinearities to minimize the PDE residuals, thereby enabling the network to learn the solution and its derivatives accordingly.
From equation (4.8), another key observation is that PirateNets can be viewed as a linear combination of basis functions at initialization. This not only allows one to control the inductive bias of the network by appropriate choice of basis, but it also enables the integration of various types of existing data into the initialization phase of the network. Specifically, suppose that there is a set of solution measurements, denoted as $\mathbf{Y}=\{y_{i}\}_{i=1}^{n}$. These can be readily used to appropriately initialize the last linear layer of the model by solving a linear least square problem,
@@
Figure 3: Physics-informed residual adaptive networks (PirateNets): In our model, input coordinates are first projected into a high-dimensional feature space using random Fourier features, then followed by passing through $N$ adaptive residual blocks. Each block consists of three dense layers, augmented with two gating operations that incorporate shallow latent features. The key design of the architecture is the adaptive skip connection with a trainable parameter $\alpha$ initialized at 00. This ensures an identity mapping for each block, thereby initializing the model as a linear combination of the coordinate embeddings. More importantly, this approach circumvents the issue of pathological initialization in deep PDE residual networks. For the final layer, we propose a physics-informed initialization based on solving a least squares problem to fit the available data, while all other weights are initialized following the Glorot scheme, and biases are set to zero. As training progresses, the model’s depth gradually increases as the nonlinearities become activated, enabling the model to progressively recover its approximation capacity.